If f(x)=2x+1 and g(x)=x^2 , find (g o f)(2).
Answer by Nate(3500)

First, define [g o f].
g(2x + 1) = (2x + 1)^2 = 4x^2 + 4x + 1
plug in 2 for x
4*4 + 4*2 + 1 = 16 + 8+ 1 = 25